OpenStudy (anonymous):

Verify the identity: 1-(sin²ß/(1-cos²ß))= -cosß

myininaya (myininaya):

\[1-\frac{\sin^2(B)}{1-\cos^2(x)}=1-\frac{\sin^2(B)}{\sin^2(B)}=1-1=0 \neq -\cos(B) \text{ for all B}\]

myininaya (myininaya):

oops that x=B by the way

OpenStudy (anonymous):

Were not sovling for ß, its verifying that 1-(sin²ß/(1-cos²ß)) does equal -cosß lol

myininaya (myininaya):

i'm not solving

myininaya (myininaya):

i showed that one side is 0

myininaya (myininaya):

-cos(B) is not 0 for all B it is sometimes 0 for some B

OpenStudy (anonymous):

OPPS ;o cos shouldn't be squared...

OpenStudy (mertsj):

In other words, you have posted the problem incorrectly,.

myininaya (myininaya):

\[1-\frac{\sin^2(B)}{1-\cos(B)}=-\cos(B)?\]

OpenStudy (anonymous):

yes

myininaya (myininaya):

i would multiply top and bottom of that fraction by the conjugate of the bottom

myininaya (myininaya):

\[1-\frac{\sin^2(B)}{1-\cos(B)} \cdot \frac{1+\cos(B)}{1+\cos(B)}=1-\frac{\sin^2(x)(1+\cos(B))}{1-\cos^2(B)}\]

myininaya (myininaya):

do you see what to do now?

myininaya (myininaya):

you can use an identity at the bottom and something will cancel

OpenStudy (anonymous):

yes, thank you! :D I have one more if you don't mind..?

myininaya (myininaya):

ok

myininaya (myininaya):

that didn't work

myininaya (myininaya):

i can't open it correctly

OpenStudy (anonymous):

oh.. one sec then

OpenStudy (mertsj):